# How (Mis)calibrated is Your Federated CLIP and What To Do About It?

### v1

@@ 4.2 Calibration error in federated learning
In this section, we compare the calibration error of TPT-based methods in both the offline and FL setting. The goal of this comparison is to understand whether the miscalibration of TPT-based methods (eg, CoOp, KgCoOp), observed by Wang et. al., [dor] in the offline CLIP fine-tuning setting, also holds true in the FL setting, which is the main focus of this work. To this end, in Tab. 2 we report the performance of three families of models: (i) TPT methods without calibration in the offline setting (eg, CoOp [coop]), (ii) TPT methods along with in-training calibration (eg, CoOp + DCA [dca]) in the offline setting, and (iii) TPT methods without calibration in the FL setting (eg, FedCoOp or PromptFL [promptfl]). The experiments were performed in the in-distribution setting, which was introduced in Sec. 4.1. Note that the offline methods operate in a centralized setting.
We observe that, on average, the TPT methods in the offline setting (first part of Tab. 2) have a worse calibration error than the same TPT methods coupled with a calibration technique (second part of Tab. 2). This is not surprising and have been empirically validated in previous works [dor]. However, what remained previously unexplored in the literature is an analysis of: fine-tuning CLIP in a distributed setting worsens calibration to what extent? In the final part of Tab. 2 we observe that the same TPT methods when employed in the FL setting, exhibit much worse calibration error across all three calibration metrics. While this may not appear as a surprise to a reader, because Chu et. al., [nucfl] have demonstrated a similar behaviour: with all else being equal, miscalibration worsens in the FL setting. However, the study in [nucfl] is limited to ResNets, where the networks have been trained from scratch in each client. In contrary, our analyses encompass fine-tuning CLIP in FL setting, which has more significant implications given the widespread adoption of CLIP in downstream tasks.
@@
The results in Tab. 2 suggest that miscalibration due to fine-tuning CLIP in FL is inevitable, as in ResNets [nucfl]. Thus, evaluating the progress of FL based on accuracy alone, as in recent state-of-the-art FL methods [fedotp, fedpha, fedpgp], gives an incomplete assessment of performance.
@@
CIFAR-10 CIFAR-100
ZS-CLIP [radford2019language] 89.91 10.16 2.62 8.54 66.84 15.06 3.84 14.98
FL setting (without calibration)
PromptFL [promptfl] 92.42 4.94 1.64 3.94 72.43 9.27 2.43 8.66
FedPGP [fedpgp] 91.28 5.89 1.78 5.17 71.13 9.91 2.59 9.78
FedOTP [fedotp] 94.38 6.03 1.74 5.37 74.81 14.44 3.63 14.30
FedPHA [fedpha] 93.72 4.59 1.47 3.52 74.23 8.91 2.33 8.61
$\text{FL}^{2}\text{oRA}$-text (Ours) 92.64 4.99 1.68 3.77 73.71 8.58 2.28 8.13
$\text{FL}^{2}\text{oRA}$-vision (Ours) 97.08 2.94 1.14 1.98 83.00 7.18 2.13 6.39
$\text{FL}^{2}\text{oRA}$ (Ours) 97.04 2.99 1.12 2.06 83.70 6.70 2.05 6.09
FL setting (with calibration: DCA [dca])
PromptFL [promptfl] 92.12 4.98 1.73 3.92 72.07 8.92 2.34 8.77
FedPGP [fedpgp] 91.54 5.65 1.75 4.69 71.36 10.38 2.73 10.14
FedOTP [fedotp] 94.30 5.41 1.59 4.53 74.63 12.67 3.27 12.52
FedPHA [fedpha] 93.53 4.81 1.46 3.41 75.30 8.06 2.20 7.63
$\text{FL}^{2}\text{oRA}$-text (Ours) 92.54 5.08 1.77 4.13 73.94 8.70 2.34 8.29
$\text{FL}^{2}\text{oRA}$-vision (Ours) 97.14 2.75 1.11 1.93 83.16 7.08 2.31 6.32
$\text{FL}^{2}\text{oRA}$ (Ours) 97.00 2.99 1.15 1.87 83.65 6.64 2.12 6.20
FL setting (with calibration: MDCA [mdca])
PromptFL [promptfl] [mdca] 92.31 5.07 1.67 3.95 72.02 9.18 2.45 8.85
FedPGP [fedpgp] 90.98 5.68 1.79 5.06 71.88 9.82 2.66 9.65
FedOTP [fedotp] 94.24 5.83 1.77 5.17 74.73 14.42 3.60 14.45
FedPHA [fedpha] 93.55 4.78 1.53 3.38 75.34 8.65 2.39 8.11
$\text{FL}^{2}\text{oRA}$-text (Ours) 92.66 5.11 1.73 3.87 73.72 8.66 2.29 8.15
$\text{FL}^{2}\text{oRA}$-vision (Ours) 97.07 2.83 1.12 1.92 82.98 7.19 2.14 6.37
$\text{FL}^{2}\text{oRA}$ (Ours) 96.92 3.03 1.15 1.96 83.73 6.82 2.01 6.13
@@ 4.3 Can we improve calibration?
Given that existing CLIP-based TPT methods worsen calibration in FL, a natural question is whether calibration can be meaningfully improved and how? To answer this, we systematically evaluate state-of-the-art TPT methods for personalized FL across three benchmarks (in-distribution, domain generalization, and base-to-new generalization), ablate two in-training calibration techniques (DCA [dca], MDCA [mdca]), and compare four global aggregation methods (FedAvg [fedavg], FedProx [fedprox], FedDyn [feddyn], FedNova [fednova]). We discuss the following:
[-] Table 3 reports results on the in-distribution benchmark using CIFAR-10 and CIFAR-100 [cifar], where classes are distributed across 100 clients with a 10% participation rate per communication round.
[-] Table 4 presents results on the domain generalization benchmark for OfficeHome [officehome], PACS [pacs], and VLCS [vlcs]. We instantiate two clients per domain and train with 100% client participation. We report performance averaged over all domains and datasets; per-domain results are in the Appendix.
[-] Finally, Table 5 evaluates the base-to-new generalization benchmark on five visual classification datasets: Food101 [food], DTD [dtd], Caltech101 [caltech], Flowers102 [oxfordflowers], and OxfordPets [oxfordpets], using 16 shots per class. We assign 10 clients per dataset with full participation. For each dataset, half of the classes are treated as base (seen) classes during training, and evaluation measures performance on both base and novel classes; we report the harmonic mean across all datasets. Detailed per-dataset results are in the Appendix.
@@
Method Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$
ZS-CLIP [radford2019language] 86.98 5.76 1.93 5.74
FL setting (without calibration)
PromptFL [promptfl] 91.25 2.88 1.21 2.79
FedPGP [fedpgp] 92.71 2.67 0.88 2.24
FedOTP [fedotp] 91.07 9.34 2.36 9.20
FedPHA [fedpha] 92.87 2.29 0.81 1.96
$\text{FL}^{2}\text{oRA}$-text (Ours) 91.49 2.84 0.95 2.58
$\text{FL}^{2}\text{oRA}$-vision (Ours) 92.93 2.23 0.84 1.98
$\text{FL}^{2}\text{oRA}$ (Ours) 93.11 2.37 0.79 1.83
FL setting (with calibration: DCA [dca])
PromptFL [promptfl] 91.32 3.07 1.22 2.87
FedPGP [fedpgp] 93.93 2.28 0.68 2.06
FedOTP [fedotp] 91.54 7.37 1.90 7.27
FedPHA [fedpha] 93.93 2.17 0.64 2.02
$\text{FL}^{2}\text{oRA}$-text (Ours) 91.40 2.63 0.97 2.47
$\text{FL}^{2}\text{oRA}$-vision (Ours) 92.89 2.56 0.86 2.05
$\text{FL}^{2}\text{oRA}$ (Ours) 93.13 2.09 0.72 1.95
FL setting (with calibration: MDCA [mdca])
PromptFL [promptfl] 91.32 2.91 1.21 2.77
FedPGP [fedpgp] 94.04 2.29 0.90 1.90
FedOTP [fedotp] 91.88 9.26 2.40 9.24
FedPHA [fedpha] 94.06 2.23 0.94 1.83
$\text{FL}^{2}\text{oRA}$-text (Ours) 91.46 2.50 0.85 2.34
$\text{FL}^{2}\text{oRA}$-vision (Ours) 92.93 2.10 0.80 1.91
$\text{FL}^{2}\text{oRA}$ (Ours) 93.07 2.24 0.77 1.64
Effect of calibration techniques. Across all federated settings, in-training calibration with DCA or MDCA yields only modest gains. As shown in Table 3, on CIFAR-10/100, applying DCA or MDCA to FL prompt-tuning methods produces small reductions in ECE and ACE for some configurations (e.g., PromptFL, FedPHA on CIFAR-100), but many methods remain noticeably miscalibrated. For instance, FedOTP on CIFAR-100 still has ECE above 12% after DCA and above 14% after MDCA. A similar pattern appears on domain generalization and base-to-new generalization (Tables 4 and 5): DCA and MDCA slightly improve FedPGP and FedPHA in some cases, yet other methods (again notably FedOTP) stay severely miscalibrated, with ECE often exceeding 7–9%. Overall, in-training calibration in TPT for FL rarely produces large, systematic improvements and sometimes even increases MCE.
Effect of global aggregation. We next test whether stronger FL optimizers can alleviate miscalibration. As shown in Tab.  6, replacing FedAvg with FedProx, FedDyn, or FedNova produces no consistent improvement in either accuracy or calibration. The ECE/MCE/ACE metrics of these optimizers closely track those of FedAvg and typically differ only within run-to-run variability; in several cases, FedAvg is slightly better. This suggests that the calibration issues of TPT for FL are not primarily driven by the choice of aggregation algorithm, but by how prompts are adapted under client heterogeneity.
@@
Method Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$
ZS-CLIP [radford2019language] 82.05 4.52 1.43 4.69
FL setting (without calibration)
PromptFL [promptfl] 82.98 4.11 1.31 3.82
FedPGP [fedpgp] 81.09 6.17 2.07 5.97
FedOTP [fedotp] 29.52 21.88 5.03 21.90
FedPHA [fedpha] 54.97 16.03 4.64 15.69
$\text{FL}^{2}\text{oRA}$-text (Ours) 86.70 3.58 1.13 3.21
$\text{FL}^{2}\text{oRA}$-vision (Ours) 84.39 4.34 1.33 4.06
$\text{FL}^{2}\text{oRA}$ (Ours) 86.25 3.91 1.41 3.20
FL setting (with calibration: DCA [dca])
PromptFL [promptfl] 84.25 3.97 1.30 3.62
FedPGP [fedpgp] 80.90 5.82 1.82 5.69
FedOTP [fedotp] 28.63 26.57 7.34 26.63
FedPHA [fedpha] 54.56 17.93 5.53 17.38
$\text{FL}^{2}\text{oRA}$-text (Ours) 86.80 3.63 1.20 2.97
$\text{FL}^{2}\text{oRA}$-vision (Ours) 84.31 3.13 1.05 3.07
$\text{FL}^{2}\text{oRA}$ (Ours) 86.08 4.21 1.61 3.32
FL setting (with calibration: MDCA [mdca])
PromptFL [promptfl] 84.62 4.06 1.23 3.73
FedPGP [fedpgp] 81.51 4.93 1.65 4.80
FedOTP [fedotp] 28.53 25.35 6.25 25.32
FedPHA [fedpha] 52.34 20.55 6.63 19.86
$\text{FL}^{2}\text{oRA}$-text (Ours) 86.32 3.42 1.07 3.11
$\text{FL}^{2}\text{oRA}$-vision (Ours) 84.37 4.17 1.31 4.00
$\text{FL}^{2}\text{oRA}$ (Ours) 86.34 4.10 1.55 3.35
Improving calibration through LoRA fine-tuning. The evidence above suggests that the core problem is not merely how we aggregate or calibrate CLIP-based TPT models, but what we fine-tune. Prompt-tuning operates only on a small set of continuous prompts while keeping CLIP’s backbone frozen. In federated settings with heterogeneous clients, this makes the prompts highly sensitive to local idiosyncrasies: each client learns sharp, overconfident decision boundaries on top of a fixed representation, and FedAvg must reconcile these conflicting prompt updates. As a result, logits are systematically overconfident and post-hoc calibration has little room to correct the mismatch between frozen features and locally adapted prompts.
A natural alternative is to adapt the backbone itself in a parameter-efficient way. Rather than relying solely on prompts to absorb all client heterogeneity, we allow the shared representation to move slightly so that the induced feature geometry is better aligned with federated label distributions. This motivates our use of LoRA-based PEFT on CLIP under the FL setting, which we denote as $\text{FL}^{2}\text{oRA}$.
@@
LoRA augments linear layers of CLIP’s vision and/or text encoders with low-rank adapters. LoRA applies to any linear transformation with weights $W\in\mathbb{R}^{m\times n}$ by expressing the update $\delta W=\alpha AB$ in low rank matrices $A\in\mathbb{R}^{m\times r},B\in\mathbb{R}^{r\times n}$, where $r$ is the LoRA rank and $\alpha$ is a fixed scaling factor (typically $\alpha=1/r$). We consider three variants of LoRA-based fine-tuning: applying LoRA to the text encoder only ($\text{FL}^{2}\text{oRA}$-text), to the vision encoder only ($\text{FL}^{2}\text{oRA}$-vision), or to both encoders ($\text{FL}^{2}\text{oRA}$). $\text{FL}^{2}\text{oRA}$, and its variants, apply LoRA to all layers of the corresponding encoder.
@@
Method FedAvg [fedavg] FedProx [fedprox]
Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$
PromptFL 72.43 9.27 2.43 8.66 73.12 8.16 2.24 7.29
FedKgCoOp 70.68 11.20 2.95 11.05 71.56 11.08 2.86 8.57
FedMaPLe 81.73 7.53 2.20 6.93 82.68 7.84 2.34 7.05
$\text{FL}^{2}\text{oRA}$-text (Ours) 73.71 8.58 2.28 8.13 74.06 8.82 2.45 8.45
$\text{FL}^{2}\text{oRA}$-vision (Ours) 83.00 7.18 2.13 6.39 82.91 7.25 2.25 6.64
$\text{FL}^{2}\text{oRA}$ (Ours) 83.70 6.70 2.05 6.09 83.55 6.86 2.14 6.15
Method FedDyn [feddyn] FedNova [fednova]
Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$
PromptFL 73.28 8.34 2.35 7.56 72.89 9.54 3.27 8.89
FedKgCoOp 71.78 11.45 2.39 7.34 70.38 12.38 3.21 8.32
FedMaPLe 82.07 7.08 2.19 7.06 81.38 8.26 2.93 8.19
$\text{FL}^{2}\text{oRA}$-text (Ours) 74.95 8.82 2.50 8.05 74.58 9.05 2.93 8.67
$\text{FL}^{2}\text{oRA}$-vision (Ours) 83.18 7.78 2.47 6.80 82.67 8.81 3.01 7.04
$\text{FL}^{2}\text{oRA}$ (Ours) 83.42 6.79 2.15 6.15 82.89 7.03 2.56 6.67
In sharp contrast to TPT, the only consistent and substantial calibration gains come from parameter-efficient fine-tuning of CLIP’s backbone, as implemented in our $\text{FL}^{2}\text{oRA}$ variants. On CIFAR-10/100 (Tab. 3), $\text{FL}^{2}\text{oRA}$-vision and $\text{FL}^{2}\text{oRA}$ achieve the best or near-best accuracy and clearly outperform all prompt-tuning baselines in ECE/MCE/ACE, both with and without calibration. For example, on CIFAR-100 without calibration, $\text{FL}^{2}\text{oRA}$ improves accuracy from 74.81% (FedOTP) to 83.70% and reduces ECE from 14.44% (FedOTP) and 8.91% (FedPHA) down to 6.70%. On the domain generalization benchmark (Tab. 4), $\text{FL}^{2}\text{oRA}$ achieves the highest accuracy without calibration (93.11%) and, after applying DCA or MDCA, offers the best ECE/ACE trade-off: $\text{FL}^{2}\text{oRA}$ attains the lowest ACE with DCA (1.95) and with MDCA (1.64), while $\text{FL}^{2}\text{oRA}$-vision achieves the best ECE and MCE under MDCA. On the more challenging base-to-new setting (Tab. 5), where all methods exhibit worse calibration and post-hoc calibration has a stronger effect, $\text{FL}^{2}\text{oRA}$-text attains the highest accuracy across all configurations, and our LoRA variants match or surpass the calibrated versions of competing baselines in both accuracy and calibration.

### v2

@@ 4.2 How does Federated Learning affect calibration?
In this section, we compare the calibration error of TPT-based methods (e.g., CoOp, KgCoOp) in both the offline and FL setting. Our goal is to understand whether the miscalibration of TPT-based methods, observed in [dor] in the offline CLIP fine-tuning setting, also holds true in the FL setting. Tab.˜2 reports the performance of TPT methods: (i) in the offline setting without calibration, (ii) in the offline setting with calibration (i.e., with DCA [dca]), and (iii) in the FL setting without calibration. In addition, we consider the visual prompt tuning method, VPT [vpt], for comprehensiveness. The FL experiments were performed in the in-distribution setting.
On average across metrics and datasets, offline prompt-tuning methods exhibit higher calibration error than when combined with a calibration technique. This aligns with prior empirical findings [dor]. A previously unexplored question is: to what extent does fine-tuning CLIP in a distributed setting worsen calibration? As shown in the last column of Tab. 2, the same methods exhibit substantially higher calibration error in FL across all metrics. While prior work [nucfl] reports similar trends, it is limited to ResNets trained from scratch. In contrast, we study CLIP fine-tuning in FL, which is particularly relevant given its widespread downstream use.
@@
In summary, the results in Tab. 2 indicate that fine-tuning CLIP in the FL setting consistently introduces miscalibration, mirroring observations for ResNets [nucfl]. Therefore, assessing CLIP-based FL methods solely based on accuracy, as done in several recent works [fedotp, fedpha, fedpgp], provides an incomplete evaluation of their performance.
@@
Method FL setting (w/o calib) FL setting (w/ calib: DCA [dca]) FL setting (w/ calib: MDCA [mdca])
Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$
CIFAR-10
PromptFL [promptfl] 92.42 4.94 1.64 3.94 92.12 4.98 1.73 3.92 92.31 5.07 1.67 3.95
FedPGP [fedpgp] 91.28 5.89 1.78 5.17 91.54 5.65 1.75 4.69 90.98 5.68 1.79 5.06
FedOTP [fedotp] 94.38 6.03 1.74 5.37 94.30 5.41 1.59 4.53 94.24 5.83 1.77 5.17
FedPHA [fedpha] 93.72 4.59 1.47 3.52 93.53 4.81 1.46 3.41 93.55 4.78 1.53 3.38
$\text{FL}^{2}\text{oRA}$ (Ours) 97.04 2.99 1.12 2.06 97.00 2.99 1.15 1.87 96.92 3.03 1.15 1.96
CIFAR-100
PromptFL [promptfl] 72.43 9.27 2.43 8.66 72.07 8.92 2.34 8.77 72.02 9.18 2.45 8.85
FedPGP [fedpgp] 71.13 9.91 2.59 9.78 71.36 10.38 2.73 10.14 71.88 9.82 2.66 9.65
FedOTP [fedotp] 74.81 14.44 3.63 14.30 74.63 12.67 3.27 12.52 74.73 14.42 3.60 14.45
FedPHA [fedpha] 74.23 8.91 2.33 8.61 75.30 8.06 2.20 7.63 75.34 8.65 2.39 8.11
$\text{FL}^{2}\text{oRA}$ (Ours) 83.70 6.70 2.05 6.09 83.65 6.64 2.12 6.20 83.73 6.82 2.01 6.13
@@ 4.3 Can we improve calibration?
Given that prompt tuning methods worsen calibration in FL, a natural question is whether calibration can be meaningfully improved and how. To answer this, we systematically evaluate state-of-the-art FL baselines in three settings: in-distribution (Tab. 3), domain generalization (Tab. 4), and base-to-new generalization (Tab. 5), with and without calibration regularization.
@@
Method FL setting (w/o calib) FL setting (w/ calib: DCA [dca]) FL setting (w/ calib: MDCA [mdca])
Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$
PromptFL [promptfl] 91.25 2.88 1.21 2.79 91.32 3.07 1.22 2.87 91.32 2.91 1.21 2.77
FedPGP [fedpgp] 92.71 2.67 0.88 2.24 93.93 2.28 0.68 2.06 94.04 2.29 0.90 1.90
FedOTP [fedotp] 91.07 9.34 2.36 9.20 91.54 7.37 1.90 7.27 91.88 9.26 2.40 9.24
FedPHA [fedpha] 92.87 2.29 0.81 1.96 93.93 2.17 0.64 2.02 94.06 2.23 0.94 1.83
$\text{FL}^{2}\text{oRA}$ (Ours) 93.11 2.37 0.79 1.83 93.13 2.09 0.72 1.95 93.07 2.24 0.77 1.64
@@
Method FL setting (w/o calib) FL setting (w/ calib: DCA [dca]) FL setting (w/ calib: MDCA [mdca])
Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$
PromptFL [promptfl] 82.98 4.11 1.31 3.82 84.25 3.97 1.30 3.62 84.62 4.06 1.23 3.73
FedPGP [fedpgp] 81.09 6.17 2.07 5.97 80.90 5.82 1.82 5.69 81.51 4.93 1.65 4.80
FedOTP [fedotp] 29.52 21.88 5.03 21.90 28.63 26.57 7.34 26.63 28.53 25.35 6.25 25.32
FedPHA [fedpha] 54.97 16.03 4.64 15.69 54.56 17.93 5.53 17.38 52.34 20.55 6.63 19.86
$\text{FL}^{2}\text{oRA}$ (Ours) 86.25 3.91 1.41 3.20 86.08 4.21 1.61 3.32 86.34 4.10 1.55 3.35
Effect of calibration techniques. A seemingly straightforward way to improve calibration is to impose calibration regularization (e.g., with DCA [dca] or MDCA [mdca]) during federated fine-tuning. We applied these techniques and report results of our experiments in Tab.˜3. Applying DCA or MDCA to FL prompt-tuning methods yields only modest improvements in ECE and ACE for certain configurations (e.g., PromptFL and FedPHA on CIFAR-100), while many methods remain substantially miscalibrated. For example, on CIFAR-100, FedOTP still exhibits ECE above 12% with DCA and above 14% with MDCA. A similar pattern appears on domain generalization and base-to-new generalization (Tabs. 4 and 5). DCA and MDCA slightly improve calibration on FedPGP and FedPHA in some cases, yet other methods (again notably FedOTP) stay severely miscalibrated, with ECE often exceeding 7–9%. Overall, we observe that in-training calibration in TPT for FL rarely produces large, systematic improvements and sometimes even makes it more miscalibrated.
Effect of global aggregation. Next, we test whether stronger FL optimizers can alleviate miscalibration. As shown in Tab.˜6, replacing FedAvg with FedProx, FedDyn or FedNova produces no consistent improvement in either accuracy or calibration. The calibration metrics of these optimizers closely track those of FedAvg and typically differ only within run-to-run variability; in several cases, FedAvg is slightly better. We conclude that calibration issues of TPT for FL are not primarily driven by the choice of global aggregation algorithm.
@@
Improving calibration through LoRA fine-tuning. Most existing FL approaches for CLIP fine-tuning rely on textual prompt tuning. To provide a different perspective, we ask: could the miscalibration issue stem from what we choose to fine-tune? To find an answer, we propose to replace text prompts with LoRA for fine-tuning CLIP. LoRA [lora] augments linear layers of CLIP’s vision and/or text encoders with low-rank adapters. LoRA applies to any linear transformation with weights $W\in\mathbb{R}^{m\times n}$ by expressing the update $\delta W=\alpha AB$ in low rank matrices $A\in\mathbb{R}^{m\times r},B\in\mathbb{R}^{r\times n}$, where $r$ is the LoRA rank and $\alpha$ is a fixed scaling factor (typically $\alpha=1/r$). We apply LoRA to the Q, K, and V projection layers of all blocks in both encoders during fine-tuning, and call our method $\text{FL}^{2}\text{oRA}$. The implementation details of $\text{FL}^{2}\text{oRA}$ are in the Appendix.
@@
Method FedAvg[fedavg] FedProx[fedprox] FedDyn[feddyn] FedNova[fednova]
Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$
PromptFL 72.43 9.27 2.43 8.66 73.12 8.16 2.24 7.29 73.28 8.34 2.35 7.56 72.89 9.54 3.27 8.89
FedKgCoOp 70.68 11.20 2.95 11.05 71.56 11.08 2.86 8.57 71.78 11.45 2.39 7.34 70.38 12.38 3.21 8.32
FedMaPLe 81.73 7.53 2.20 6.93 82.68 7.84 2.34 7.05 82.07 7.08 2.19 7.06 81.38 8.26 2.93 8.19
$\text{FL}^{2}\text{oRA}$ (Ours) 83.70 6.70 2.05 6.09 83.55 6.86 2.14 6.15 83.42 6.79 2.15 6.15 82.89 7.03 2.56 6.67
From Tab. 3, we observe that $\text{FL}^{2}\text{oRA}$ clearly outperforms all the prompt-tuning baselines in all three ECE/MCE/ACE metrics, even without needing additional regularization. For example, on CIFAR-100 without calibration, $\text{FL}^{2}\text{oRA}$ improves accuracy from 74.81% (FedOTP) to 83.70% and reduces ECE from 14.44% (FedOTP) and 8.91% (FedPHA) down to 6.70%. On the more challenging base-to-new setting (Tab. 5), where all methods exhibit worse calibration and additional calibration has some effect, $\text{FL}^{2}\text{oRA}$ demonstrates all-round performance across all configurations. On the domain generalization benchmark (Tab. 4), the gap between prompt-tuning methods and $\text{FL}^{2}\text{oRA}$ becomes smaller, as DG poses a less severe adaptation challenge for CLIP: the input resolution and the label set remain unchanged, and only the image distribution shifts. Nevertheless, $\text{FL}^{2}\text{oRA}$ still achieves the highest accuracy without calibration (93.11%) and, after applying DCA or MDCA, provides the best calibration trade-off, attaining the lowest ACE with DCA (1.95) and MDCA (1.64).